# CAMAR: Continuous Actions Multi-Agent Routing

### v1

@@ Appendix A Appendix A — Evaluation Protocol
  • Medium trains on one map setting but evaluates on 12 variants that share the same map types while changing agent count and obstacle density. This probes generalisation within the same domain and includes testing generalization across agents.
@@ Appendix E Appendix E — Extended Scalability Analysis
We also study CAMAR scalability in comparison to other popular MARL environments to better understand its limits.
@@ Appendix F Appendix F — Future Work
We highlight several research directions that can be explored using CAMAR.
First, CAMAR allows testing with many agents, which supports the development of scalable multi-agent reinforcement learning (MARL) methods. Future work can focus on improving how algorithms handle hundreds of agents in complex environments.
Second, CAMAR is a good platform for studying communication between agents. In large-scale tasks, agents often need to share information with nearby teammates. This is an important challenge that is still not well studied. CAMAR provides the tools to explore this type of localized communication.
Third, CAMAR includes realistic navigation tasks, which makes it useful for combining planning methods (like RRT, RRT*) with learning-based agents. This combination could improve both sample efficiency and robustness by using long-term planning together with learned policies.
Fourth, CAMAR supports agents with different abilities and dynamics. This enables research on heterogeneous teams, where agents may have different sizes, speeds, or types of control. In future work, researchers can develop methods that automatically adapt to such diverse agent groups.
@@ Appendix H Appendix H — Extended Related Work
The Waterworld environment, part of the SISL (Stanford Intelligent Systems Laboratory) suite  [24], is a continuous control benchmark where multiple agents move in a bounded two-dimensional space to collect moving targets (“food”) while avoiding harmful objects (“poison”). Agents have continuous observations and actions, and their motion dynamics are simple and lightweight, which allows high simulation speeds for training.
@@
The Multi-Robot Warehouse environment [11] simulates robots moving in a warehouse to collect and deliver requested goods. The layout is grid-based, and agents must coordinate to navigate around shelves and other robots. The default version is implemented in Python using a discrete state and action space. It supports partial observability but does not provide procedural generation: the warehouse layout is fixed. As a result, agents do not need to generalize across different maps.
@@
The Trash Pickup environment [25] is a grid-based multi-agent task where agents move around a map to collect pieces of trash and deliver them to designated drop-off locations. It is implemented in C with a Python API for controlling agents, which allows efficient simulation. The environment uses discrete state and action spaces and supports partial observability.
@@
The StarCraft Multi-Agent Challenge [18] is one of the most widely used benchmarks in the MARL community. In SMAC, a team of StarCraft II units controlled by independent agents must cooperate to defeat an opposing team. The environment is partially observable, and by default uses discrete action spaces. The underlying maps are fixed, which allows agents to solve the tasks without relying on observation inputs by simply memorizing the optimal action sequence for each map. This significantly reduces the need for generalization across scenarios.
@@
SMACv2 [15] addresses one of the key limitations of SMAC by introducing randomly generated maps with random positions of units, which prevents memorization of fixed action sequences and forces agents to generalize their policies across start positions. The rest of the environment remains the same as SMAC, with partially observable states and discrete actions by default. Like SMAC, SMACv2 does not include an evaluation protocol, but it can be evaluated using the protocol proposed in [32].
@@
SMAX [17] is a JAX-based re-implementation of SMACv2 that leverages hardware acceleration for faster simulation. In addition to performance improvements, SMAX introduces the option to switch from discrete to continuous action spaces, making it more flexible for testing different types of MARL algorithms.
@@
The Multi-Agent Particle Environment [20, 16] is a lightweight 2D simulator for testing cooperative, competitive, and communication multi-agent tasks. Agents and landmarks are represented as simple geometric shapes, and their interactions follow basic physical dynamics. MPE supports both continuous and discrete action spaces, and scenarios can be either fully or partially observable. Procedural generation is not used in the default scenarios, although obstacle or landmark positions can be randomized in some tasks. This means generalization across different maps is not strictly required. The environment does not include a built-in evaluation protocol.
@@
The JaxMARL re-implementation of MPE [17] offers hardware acceleration through JAX, which enables much faster simulation compared to the original Python implementation. Functionally, it mirrors MPE in terms of available scenarios, physics, and agent capabilities. It supports both continuous and discrete action spaces. Like the original MPE, it does not include procedural map generation by default and does not provide an evaluation protocol.
@@
JaxNav [26] is a navigation-focused benchmark implemented within the JaxMARL framework. It places agents in a continuous 2D space where they must navigate to goal locations while avoiding static obstacles. The environment supports both continuous and discrete action spaces. Maps are fixed in the default setup, so agents do not need to generalize to unseen layouts. JaxNav is implemented in JAX for hardware acceleration but does not provide an evaluation protocol.
@@
Nocturne [27] is a 2D partially observed driving simulator implemented in C++ for performance, with a Python API for training and evaluation. It focuses on realistic autonomous driving scenarios sourced from the Waymo Open Dataset. The environment provides a set of benchmark tasks in which agents control autonomous vehicles interacting with traffic participants whose trajectories are taken from recorded data. This setup captures complex multi-agent interactions requiring coordination, prediction of other agents’ intentions, and handling of partial observability.
@@
The Partially Observable Grid Environment for Multiple Agents [12] is a grid-based benchmark for partially observable multi-agent pathfinding. Agents operate with ego-centric local observations and must reach individual goals while avoiding collisions. A key feature of POGEMA is its procedural map generation, producing diverse layouts—random maps, mazes, and warehouse-style scenarios—that require agents to generalize to unseen environments. It also offers integrated evaluation protocols with standardized metrics, enabling consistent comparison across reinforcement learning, planning, and hybrid methods. Implemented in Python, POGEMA supports scalable multi-agent experiments with both classical and learning-based policies.
@@
The Vectorized Multi-Agent Simulator for Collective Robot Learning [14] is a PyTorch-based, vectorized 2D physics framework designed for efficient multi-robot benchmarking. It includes a modular interface for defining custom scenarios, alongside a set of built-in multi-robot tasks, that each involve a relatively small number of agents. VMAS stands out for its GPU-accelerated, batch-simulated environments. It supports inter-agent communication and sensors (e.g., LIDAR).
@@
Scalable Multi-Agent Realistic Testbed [28] is a physics-based simulator tailored for bridging Multi-Agent Path Finding (MAPF) algorithms and real-world performance. Designed for scalability, SMART supports simulation of thousands of agents, enabling evaluation of large-scale deployments. However, its simulation speed is relatively low (around 335 SPS for 100 agents), which can limit large-scale reinforcement learning experiments. The platform targets both academic researchers and industrial users who lack access to extensive physical robot fleets, offering a realistic environment to test MAPF performance under near-real conditions.
@@
Gazebo is a widely used open-source 3D robotics simulator that integrates tightly with ROS. It supports detailed physics simulation, realistic sensors (e.g., LIDAR, cameras), and high-quality rendering. While Gazebo allows multi-robot experimentation and moderate scalability, it does not support GPU acceleration for simulation logic - GPU use is limited to rendering and sensor visuals only. This limitation can make simulations of large multi-agent systems slow or resource-heavy.
@@
Webots is a commercial-grade 3D robotics simulator offering a wide array of built-in robot models, sensors, and actuator libraries. It provides realistic physics simulation, high-quality graphics, and ROS support. Webots is well-suited for prototyping and academic or industrial robotic research. However, from a multi-agent perspective, it’s not optimized for running large numbers of agents at scale—simulation speed tends to drop significantly as agent count grows, making massive multi-robot evaluations challenging.
@@
ARGoS [23] is an open-source, modular simulator designed specifically for swarm robotics and large-scale multi-agent systems. Its architecture supports running thousands of simple agents efficiently by combining multiple physics engines and leveraging parallel computation. ARGoS is customizable, allowing researchers to define their own robots, sensors, actuators, etc.

### v2

@@ Appendix A Appendix A — Evaluation Protocol
  • Medium trains on one map setting but evaluates on 12 variants that share the same map types while changing agent count and obstacle density. This probes generalization within the same domain and includes testing generalization across agents. The current protocol is designed for reproducibility and fair comparison across methods. However, since the environment scales efficiently, we plan to extend the Medium tier in future releases to include a wider range of agent counts and task complexities as MARL methods progress.
@@ Appendix D Appendix D — Extended Benchmark
Detailed Benchmark Analysis
The main evaluation compares success rate (SR), flowtime (FT), makespan (MS), and coordination (CO) on two map types. On random_grid, MAPPO gives the highest SR with strong CO. RRT*+MAPPO reaches similar SR but with lower FT and MS, which shows that planning improves movement efficiency. RRT*+IPPO gives smaller gains but still improves over IPPO. The classic RRT*+PD baseline reaches high SR but low CO because it plans for each agent alone and does not react to dynamic agent movement.
Hybrid methods with off-policy algorithms show mixed behavior. RRT* improves ISAC and IDDPG in SR and in some cases CO. But RRT* weakens MASAC and MADDPG, where SR is almost zero. These algorithms use a centralized critic with long input vectors, which include observations and RRT* samples from all agents. This leads to unstable training.
IPPO and IDDPG reach good SR but slower FT than MAPPO. ISAC gives high CO but low SR. MASAC and MADDPG fail to explore and reach very low SR. Additional experiments and analysis on DDPG and SAC can be found below.
The labmaze_grid map is harder because of sparse rewards and narrow corridors. All MARL baselines drop in SR. RRT*+PD performs best in SR due to full-path planning, but CO is still low. The simple RRT+PD baseline gives worse paths and lower SR than RRT*+PD, but in some cases it still outperforms MARL baselines in SR.
These results show a clear trade-off. Planning gives strong reliability and better path efficiency. MARL gives better coordination but fails in sparse reward maps. Hybrid methods give only small gains. One reason is that the current policy networks are small and cannot process the large inputs that combine local observations with RRT* samples. Stronger model architectures may improve this integration.
@@ Appendix E Appendix E — Extended Scalability Analysis
CAMAR vs VMAS
This section provides visualizations of scenarios of CAMAR and VMAS. Fig. 7(a) and Fig. 14 use the same map generation settings for both CAMAR and VMAS to allow a fair comparison under matched conditions.
(a) CAMAR
(b) VMAS
@@ Appendix F Appendix F — Future Work
We highlight several research directions that can be explored using CAMAR:
  1. 1.
    CAMAR enables testing with large agent populations, supporting the development of scalable multi-agent reinforcement learning (MARL) methods. Future work can focus on improving how algorithms handle hundreds of agents operating in dense, dynamic environments.
  2. 2.
    CAMAR provides a natural platform for studying inter-agent communication. Large-scale coordination often requires agents to exchange information with nearby teammates, which remains an open challenge. The environment offers the necessary tools to explore such localized communication mechanisms.
  3. 3.
    CAMAR includes realistic continuous navigation tasks, making it suitable for integrating classical planning approaches (RRT, RRT) with learning-based agents. Combining long-term planning with reinforcement learning could improve both sample efficiency and robustness.
  4. 4.
    CAMAR supports agents with different dynamics and properties, enabling research on heterogeneous teams. We plan to extend the current list of dynamics. For the moment, we include differential-drive and holonomic robots, but we are going to support more options, such as car-like and 2D quadrotor dynamics.
  5. 5.
    CAMAR’s fully modular design allows users to easily redefine observation and reward functions through wrapping. We plan to introduce additional observation types, including simulated 2D LiDAR, to further diversify sensing modalities for navigation tasks.
  6. 6.
    Another direction of future extensions is to extend classical baselines beyond the current RRT+PD and RRT setup by adding widely used multi-robot navigation methods such as PRM+PD (Probabilistic roadmap for global planning with PD controller), ORCA, and MPPI. Similarly, we will expand learning-based baselines to include RNN-based algorithms and architectures using graph attention or transformers, going beyond the standard MLP-based backbones.
  7. 7.
    We aim to incorporate post-generation connectivity checks to ensure that all agents can reach their goals. While computationally expensive, this process will improve the reliability and fairness of evaluation.
  8. 8.
    CAMAR can be extended beyond multi-agent pathfinding. Future releases will include new task types such as pick-up-and-delivery, cooperative transportation, and dynamic obstacle scenarios, making CAMAR more general testbed for large-scale MARL research.
@@ Appendix H Appendix H — Extended Related Work
To sum up, our benchmark builds on the strengths of existing environments and aims to complement ongoing efforts in MARL research. Over the years, a range of benchmarks has helped researchers tackle challenges in coordination, planning, and navigation. Well-known examples include SMAC [20, 17, 19], Jumanji [21], POGEMA [14, 15], MPE [22, 18], and VMAS, each of which introduces useful features for different kinds of tasks.
SMAC [20, 17, 19] is popular for testing strategic decision-making, but it uses discrete actions and does not scale well for large environments. Jumanji [21] supports GPU acceleration and procedural generation, but its focus is not on navigation or planning. POGEMA [14, 15] is strong in grid-based navigation and procedural tasks with many agents, but it does not use continuous states or actions, which are important in robotics.
MPE [22, 18] has played an important role in early MARL research, but it cannot scale to hundreds of agents efficiently. VMAS [16] builds on MPE [22, 18] by adding its physics and continuous dynamics, making it better for robotics. However, it can still be slow and difficult to scale to larger agent populations or complex maps.
Waterworld (SISL) [34] link
RWare [13] link
RWare (Jumanji) [21] link
RWare (Pufferlib) [35] link
Trash Pickup (Pufferlib) [35] link
SMAC [20] link
SMACv2 [17] link
SMAX (JaxMARL) [19] link
[…]
JaxNav (JaxMARL) [36] link
Nocturne [37] link
POGEMA [14] link
[…]
SMART  [38] link
Gazebo  [23] link
Webots  [24] link
ARGoS  [25] link
CAMAR (Ours) link
Finally, popular robotics simulators such as Gazebo [23], Webots [24], and ARGoS [25] are powerful tools for realistic physics-based simulation and robotic deployment. Each of these frameworks serves specific use cases focused on accurate physical modeling and hardware-oriented evaluation. In contrast, our environment, CAMAR, is designed to complement these tools by focusing on large-scale MARL and benchmarking, particularly for continuous MAPF tasks. Rather than replacing existing simulators, CAMAR extends the set of available research tools, bridging the gap between large-scale MARL studies and robotics-oriented simulation.
We compare MARL environments and robotic simulators using the following criteria (Tables 16):
Continuous Observations and Actions
Robots usually operate in continuous state and action spaces, so it’s important for benchmarks to reflect these conditions. Environments like VMAS and MPE [22, 18] use continuous actions and observations, while others, like Trash Pickup [35] or POGEMA [15], use discrete representations, which limit realism in robotics simulations.
GPU Support
In multi-agent environments, each agent often has its observations and rewards. This can lead to heavy data transfers between the CPU and the GPU, especially when training deep RL models. Benchmarks like Jumanji [21] and SMAX [19] are initially built with GPU acceleration, helping to reduce training time. In contrast, many older environments like SMAC [20, 17] and MPE [22, 18] do not support GPU-based simulation.
Scalability >500 Agents
When the number of agents grows, decision-making becomes harder. A good benchmark should scale well to hundreds or thousands of agents. Environments like POGEMA [14, 15] and Trash Pickup [35] can run with thousands and even millions of agents. Others, like VMAS, MPE [22, 18], and SMAC [20, 17], are limited to much smaller groups.
Partially Observable
In most real-world scenarios, agents see only part of the environment. This is known as partial observability and is a common feature in almost every RL environment. SMAC [20, 17], RWare [13], and POGEMA [14, 15] support this feature by limiting each agent’s view.
Heterogeneous Agents
In real life, agents and robots can have different sensors, shapes, or goals. Benchmarks like VMAS and SMAC [20, 17] allow agents to behave differently, making cooperation more complex. Others, like POGEMA [14, 15] or RWare [13], usually involve homogeneous agents.
Performance >10K Steps/s
High simulation speed is essential when agents need millions of interactions to learn. For example, the Pufferlib [35] environment can run with a speed of up to 1M steps per second. This allows researchers to train and develop models more quickly. Environments like SMAC [20, 17] or VMAS can be slower, especially when running with many agents.
Python Based
Using Python makes it easier for researchers to understand and modify the environment. SMAX [19], VMAS, and POGEMA [14, 15] are implemented entirely in Python. This helps with faster development and integration into learning frameworks.
Procedural Generation
Procedural generation helps create diverse tasks that reduce the chance of overfitting. POGEMA [14, 15] and Jumanji [21] use this method to generate complex tasks automatically. Other environments, like RWare [13], use fixed layouts that limit diversity.
Requires Generalization
Only a few environments offer separate training and testing tasks, making it difficult to test whether agents generalize well to new situations. POGEMA [14, 15] and SMACv2 [17] include test scenarios that allow researchers to check generalization. Many others use the same tasks during both training and testing.
Evaluation Protocols
To compare algorithms fairly, we need well-defined test cases and metrics. Benchmarks like SMACv2 [17] and Jumanji [21] include evaluation protocols to make results reproducible and meaningful. Many other environments do not have standard evaluation tools or test setups.
Tests & CI
PyPi Listed
When an environment is available on PyPi, it becomes easier for others to install and use. Benchmarks like Jumanji [21], and Trash Pickup [35] are available on PyPi. This lowers the barrier to entry and helps increase adoption in the research community.
Here are detailed descriptions and analysis of each environment used for the comparison in the Tables 16:
@@
— The Waterworld environment, part of the SISL (Stanford Intelligent Systems Laboratory) suite  [34], is a continuous control benchmark where multiple agents move in a bounded two-dimensional space to collect moving targets (“food”) while avoiding harmful objects (“poison”). Agents have continuous observations and actions, and their motion dynamics are simple and lightweight, which allows high simulation speeds for training.
@@
— The Multi-Robot Warehouse environment [13] simulates robots moving in a warehouse to collect and deliver requested goods. The layout is grid-based, and agents must coordinate to navigate around shelves and other robots. The default version is implemented in Python using a discrete state and action space. It supports partial observability but does not provide procedural generation: the warehouse layout is fixed. As a result, agents do not need to generalize across different maps.
@@
— The Trash Pickup environment [35] is a grid-based multi-agent task where agents move around a map to collect pieces of trash and deliver them to designated drop-off locations. It is implemented in C with a Python API for controlling agents, which allows efficient simulation. The environment uses discrete state and action spaces and supports partial observability.
@@
— The StarCraft Multi-Agent Challenge [20] is one of the most widely used benchmarks in the MARL community. In SMAC, a team of StarCraft II units controlled by independent agents must cooperate to defeat an opposing team. The environment is partially observable, and by default uses discrete action spaces. The underlying maps are fixed, which allows agents to solve the tasks without relying on observation inputs by simply memorizing the optimal action sequence for each map. This significantly reduces the need for generalization across scenarios.
@@
— SMACv2 [17] addresses one of the key limitations of SMAC by introducing randomly generated maps with random positions of units, which prevents memorization of fixed action sequences and forces agents to generalize their policies across start positions. The rest of the environment remains the same as SMAC, with partially observable states and discrete actions by default. Like SMAC, SMACv2 does not include an evaluation protocol, but it can be evaluated using the protocol proposed in [29].
@@
— SMAX [19] is a JAX-based re-implementation of SMACv2 that leverages hardware acceleration for faster simulation. In addition to performance improvements, SMAX introduces the option to switch from discrete to continuous action spaces, making it more flexible for testing different types of MARL algorithms.
@@
— The Multi-Agent Particle Environment [22, 18] is a lightweight 2D simulator for testing cooperative, competitive, and communication multi-agent tasks. Agents and landmarks are represented as simple geometric shapes, and their interactions follow basic physical dynamics. MPE supports both continuous and discrete action spaces, and scenarios can be either fully or partially observable. Procedural generation is not used in the default scenarios, although obstacle or landmark positions can be randomized in some tasks. This means generalization across different maps is not strictly required. The environment does not include a built-in evaluation protocol.
@@
— The JaxMARL re-implementation of MPE [19] offers hardware acceleration through JAX, which enables much faster simulation compared to the original Python implementation. Functionally, it mirrors MPE in terms of available scenarios, physics, and agent capabilities. It supports both continuous and discrete action spaces. Like the original MPE, it does not include procedural map generation by default and does not provide an evaluation protocol.
@@
— JaxNav [36] is a navigation-focused benchmark implemented within the JaxMARL framework. It places agents in a continuous 2D space where they must navigate to goal locations while avoiding static obstacles. The environment supports both continuous and discrete action spaces. Maps are fixed in the default setup, so agents do not need to generalize to unseen layouts. JaxNav is implemented in JAX for hardware acceleration but does not provide an evaluation protocol.
@@
— Nocturne [37] is a 2D partially observed driving simulator implemented in C++ for performance, with a Python API for training and evaluation. It focuses on realistic autonomous driving scenarios sourced from the Waymo Open Dataset. The environment provides a set of benchmark tasks in which agents control autonomous vehicles interacting with traffic participants whose trajectories are taken from recorded data. This setup captures complex multi-agent interactions requiring coordination, prediction of other agents’ intentions, and handling of partial observability.
@@
— The Partially Observable Grid Environment for Multiple Agents [14] is a grid-based benchmark for partially observable multi-agent pathfinding. Agents operate with ego-centric local observations and must reach individual goals while avoiding collisions. A key feature of POGEMA is its procedural map generation, producing diverse layouts—random maps, mazes, and warehouse-style scenarios—that require agents to generalize to unseen environments. It also offers integrated evaluation protocols with standardized metrics, enabling consistent comparison across reinforcement learning, planning, and hybrid methods. Implemented in Python, POGEMA supports scalable multi-agent experiments with both classical and learning-based policies.
@@
— The Vectorized Multi-Agent Simulator for Collective Robot Learning [16] is a PyTorch-based, vectorized 2D physics framework designed for efficient multi-robot benchmarking. It includes a modular interface for defining custom scenarios, alongside a set of built-in multi-robot tasks, that each involve a relatively small number of agents. VMAS stands out for its GPU-accelerated, batch-simulated environments. It supports inter-agent communication and sensors (e.g., LIDAR).
@@
— Scalable Multi-Agent Realistic Testbed [38] is a physics-based simulator tailored for bridging Multi-Agent Path Finding (MAPF) algorithms and real-world performance. Designed for scalability, SMART supports simulation of thousands of agents, enabling evaluation of large-scale deployments. However, its simulation speed is relatively low (around 335 SPS for 100 agents), which can limit large-scale reinforcement learning experiments. The platform targets both academic researchers and industrial users who lack access to extensive physical robot fleets, offering a realistic environment to test MAPF performance under near-real conditions.
@@
— Gazebo is a widely used open-source 3D robotics simulator that integrates tightly with ROS. It supports detailed physics simulation, realistic sensors (e.g., LIDAR, cameras), and high-quality rendering. While Gazebo allows multi-robot experimentation and moderate scalability, it does not support GPU acceleration for simulation logic - GPU use is limited to rendering and sensor visuals only. This limitation can make simulations of large multi-agent systems slow or resource-heavy.
@@
— Webots is a commercial-grade 3D robotics simulator offering a wide array of built-in robot models, sensors, and actuator libraries. It provides realistic physics simulation, high-quality graphics, and ROS support. Webots is well-suited for prototyping and academic or industrial robotic research. However, from a multi-agent perspective, it’s not optimized for running large numbers of agents at scale—simulation speed tends to drop significantly as agent count grows, making massive multi-robot evaluations challenging.
@@
— ARGoS [25] is an open-source, modular simulator designed specifically for swarm robotics and large-scale multi-agent systems. Its architecture supports running thousands of simple agents efficiently by combining multiple physics engines and leveraging parallel computation. ARGoS is customizable, allowing researchers to define their own robots, sensors, actuators, etc.